Tim can fill the pool in 30 minutes, Sally fills pool in 45 minutes and bob fills the pool in 90 minutes. How long will it take for them to fill the pool if they work together?
Answer by edjones(8007)

Tim does 1/30 of the job in 1 minute.
Sally does 1/45 of the job in 1 min.
Bob does 1/90 of the job in 1 min.
1/30 + 1/45 + 1/90 =
3/90 + 2/90 + 1/90 = 6/90
=1/15 of the job done by the 3 of them in 1 min.
15 min for the 3 of them to do the job.
.
Ed